OpenStudy (anonymous):

Can you pls gimme all the equations for electric current and circuit chapter.....thnx

OpenStudy (anonymous):

@lolo123... @PFEH.1999

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@nspire

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@abhisar

OpenStudy (anonymous):

ok,wait

OpenStudy (anonymous):

well,it's impossible for us to give u all the equations!!!!

OpenStudy (anonymous):

okay only the main ones plss

OpenStudy (anonymous):

sir like only the main equations like work done formulae

OpenStudy (anonymous):

ok just give me an equation you're looking for , because i want to know that you're looking for what kinds of them...

OpenStudy (anonymous):

P.D

OpenStudy (anonymous):

and also W.D (work done)

OpenStudy (anonymous):

you want electric formulas or work/power and etc formulas?

OpenStudy (anonymous):

electric

OpenStudy (anonymous):

so plz give me an example of whatever you're looking for in electric.

OpenStudy (anonymous):

i don't know what formulas do you want about electrics.

OpenStudy (anonymous):

like formulae for finidng currrent

OpenStudy (anonymous):

well,for example R = V / I ?

OpenStudy (anonymous):

or I = q / t?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ok these are good to: U = I^2 . R . t P = I^2 . R . t / t = I^2 . R

OpenStudy (anonymous):

and i only remember these, are u looking for something else?

OpenStudy (anonymous):

thnx alot

OpenStudy (anonymous):

you're welcome ;)

OpenStudy (anonymous):

u knwo any amsdin or mod who is online

OpenStudy (anonymous):

one of my friends http://openstudy.com/users/ganeshie8

OpenStudy (anonymous):

he's awesome in mathematics, but i think he's good in physics to.

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

@dum-dum , u don't have any question?

OpenStudy (anonymous):

u got your answers?

OpenStudy (anonymous):

ya gee thnx

OpenStudy (anonymous):

you're welcome; bye :) ;)

OpenStudy (anonymous):

just please close the topic

OpenStudy (radar):

A very good law is Ohm's law which gives you I = E/R. Power P = R*I^2 for AC & DC AC power = EI times the power factor (cosine of the phase angle between current and voltage) For AC Impedance \[Z=\sqrt{R ^{2}+ X ^{2}}\]Where X = reactance. The resistor color code should be memorized if you are going to actually be working with electronics

OpenStudy (sachin00007):

V=IR Q=IT E=Q*V V=voltage F=Q*v v=velocity P=IV P=I^2*R P=V^2*R

OpenStudy (radar):

The last post has an error, the last one P=V^2*R is incorrect. P=IV, but I=V/R, so P = V*V/R which is P = V^2/R